OpenStudy (anonymous):

How many asymptotes would you expect after analyzing the equation? Select one: a. 0 Asymptotes b. 1 Vertical Asymptote c. 1 Horizontal Asymptote d. 1 Vertical and 1 Horizontal Asymptote

OpenStudy (anonymous):

man you are keeping me busy!!

OpenStudy (anonymous):

wouldnt there be 2 vertical and 1 horizontal

OpenStudy (anonymous):

hahah thannnnkkkk youuuuu

OpenStudy (anonymous):

this is the same equation as the last one right?

OpenStudy (anonymous):

im keeping that smartscore up lol

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yeah right in any case we gotta factor and cancel

OpenStudy (anonymous):

vertical x=1, x=-2 horizontal y=-1 right

OpenStudy (anonymous):

\[\frac{3-2x-x^2}{x^2+x-2}=\frac{(3-x)(1+x)}{x+1)(x+2)}=-\frac{x+3}{x+2}\]

OpenStudy (anonymous):

no it looks like it but not after you factor and cancel

OpenStudy (anonymous):

only one vertical \(x=-2\)

OpenStudy (anonymous):

there is also the horizontal one you said \(y=-1\)

OpenStudy (anonymous):

go with D

OpenStudy (anonymous):

ok